OpenStudy (amistre64):

Is 1/2 a standard when trying to determine an error approximation?

OpenStudy (amistre64):

for example: how many terms of the series: 1- 1/2 + 1/3 - 1/4 + ... are needed to compute the sum correct to two decimal places?

OpenStudy (amistre64):

the answer shows: \[a_n=1/n\text{ so choose "n" such that...}\] \[\frac{1}{n+1}<\frac{1}{2}x10^{-2}\] why the 1/2?

OpenStudy (anonymous):

Are you going up to infiniti \[\sum _{n=1}^{\infty } (-1)^{n+1}\left(\frac{1}{n}\right)\]

OpenStudy (anonymous):

?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Okay, it converges to ln(2)

OpenStudy (amistre64):

right, the sum of the series is ln(2) but is there any reason why they used ".005" in their error approximation to 2 decimal points? Why not use .001 or .007 or someelse?

OpenStudy (anonymous):

I know there are better reason, but if there are none it might be because it is in middle